# On the Influence of Pluto on Twotino Dynamics Through Their Mutual 4:3 Mean Motion Resonance

### v1

@@ Abstract
Methods. We run high-resolution, 10 Myr REBOUND simulations of the observed Twotino population in the Kuiper belt, under the perturbations of the Sun, the four giant planets, and Pluto, as massive objects.
Results. We find that all objects trapped in the 2:1 MMR with Neptune are locked in a weak 4:3 MMR with Pluto. The 4:3 resonant argument of most objects trapped in the leading and trailing islands of the 2:1 MMR, librates with amplitudes lower than $360^{\circ}$. Objects in the symmetric islands of the 2:1 MMR librate in the 4:3 MMR with amplitudes greater than $360^{\circ}$, but, contrary to circulating objects, will oscillate by up to $840^{\circ}$ visiting preferred regions on Pluto’s co-rotating frame, indicating a diluted resonant effect that may also perturb their orbits on secular timescales.
Conclusions. The importance of Pluto in shaping the structure of the trans-Neptunian region should be reconsidered, especially for resonant populations. Moreover, with current computational power, its exclusion from simulations cannot be justified.
@@ 1 Introduction
Long after Clyde Tombaugh’s discovery of Pluto in 1930, the first long-term integration of its orbit (Cohen and Hubbard, 1965), aided by the first computers, enabled the determination of its resonant state with Neptune.
Some of the pioneering works that studied the secular dynamics of the outer Solar System, from the late 1980s to the early 1990s, considered the perturbations of the Sun and the four giant planets, together with larger and larger numbers of test particles (see, e.g. Duncan et al., 1988; Torbett, 1989; Gladman and Duncan, 1990; Levison and Duncan, 1993). It probably seemed wasteful to include other massive bodies, since Uranus is over 6,500 times more massive than Pluto [^1] See e.g., https://ssd.jpl.nasa.gov/planets/phys_par.html, and references therein., and over 5,200 more massive than Eris, the two most massive currently known trans-Neptunian objects (TNO).
It has also been recognized that Pluto has an important influence over objects in its vicinity, being responsible, among other things, for the destabilization of other Plutinos, which contribute to the injection rate of comets into the inner Solar System (Morbidelli, 1997; Yu and Tremaine, 1999; Nesvorný et al., 2000; Tiscareno and Malhotra, 2009; Muñoz-Gutiérrez et al., 2019). However, Pluto can have a secular effect on other objects besides its neighboring Plutinos. In particular, Twotinos in 2:1 mean motion resonance (MMR) with Neptune appear to be directly affected by Pluto, exhibiting a significantly larger leaking rate from the resonance when Pluto’s mass is considered compared to when it is ignored (Muñoz-Gutiérrez et al., 2025).
Although Pluto is not massive enough to capture other TNOs into its own MMRs, Neptune has captured enough objects in its resonances, and those will therefore be at the same time in resonance with Pluto. In this letter, we will show that Pluto’s effect on Twotinos may be due to a previously unrecognized 4:3 MMR between them. This makes Pluto an important ingredient when studying the secular evolution of the outer Solar System, and with current computational capabilities, its exclusion from simulations cannot be justified.
This letter is organized as follows. Section 2 presents the connection of Pluto and Twotinos with Neptune. Section 3 provides the validation of the 4:3 MMR in numerical simulations. Section 4 discusses the dynamical significance of the 4:3 resonance and the effect of Eris. Finally, Section 5 presents our conclusions.
@@ 2 The Connection Through Neptune
Pluto is locked in a 3:2 MMR with Neptune, and by definition, a Twotino is locked in a 2:1 MMR with Neptune. As they are both locked in resonance with Neptune, indirectly, they are also locked in resonance with each other. While Neptune orbits six times around the Sun, Pluto will orbit four times, and a Twotino will orbit three times. Therefore, Twotinos will be in a 4:3 MMR with Pluto.
Figure 1: Evolution of the 2:1 resonant arguments with respect to Neptune (top panel) and the 4:3 resonant arguments with respect to Pluto (bottom panel) for four selected objects. The blue ($2015\,{\rm VH}_{166}$, leading), green ($2015\,{\rm GZ}_{54}$, trailing), and red ($2015\,{\rm KX}_{173}$, symmetric) curves show the behavior in the three possible Twotino libration islands. The gray curve shows a circulating ($1999\,{\rm RX}_{215}$, non-resonant) object. The vertical dashed line is placed at $62,000$ yr, representing one libration period of the symmetric object.
It is well known that Pluto’s resonant argument in the 3:2 MMR is given by:
@@
while Twotinos show libration of the angle:
$\phi_{2:1}=2\lambda_{\rm T}-\lambda_{\rm N}-\varpi_{\rm T},$ (2)
where $\lambda$ is the mean longitude, $\varpi$ the longitude of perihelion, and the sub-indices P, N, and T stand for Pluto, Neptune, and Twotinos, respectively. To cancel the dependence on Neptune’s mean longitude, we can subtract $\phi_{3:2}$ from twice the values of $\phi_{2:1}$.
@@
The libration of such resonant argument would correspond to Twotinos being locked in a first order, 4:3 MMR with Pluto (with $\phi_{4:3}$ corresponding to eq. 4D1.4 in Murray and Dermott, 1999). This resonance is weaker than the ones involving Neptune, as its relative strength includes a third-order eccentricity term ($e_{P}e_{T}^{2}$), instead of first-order terms as other resonant arguments with Neptune do ($e_{P}$, and $e_{T}$, respectively); it is also slightly more complex as it involves the longitudes of perihelion of both resonant objects.
@@
The libration of Pluto’s resonant argument with respect to Neptune has a large amplitude (from $\sim 160^{\circ}$ to $\sim 172^{\circ}$, e.g. Milani et al., 1989; Malhotra and Ito, 2022). There are also relatively large librations of Twotinos in the three possible resonant islands relative to Neptune, especially for Twotinos in the symmetric island (e.g. Figure 4 of Muñoz-Gutiérrez et al., 2025); therefore, there will be some libration between Pluto and Twotinos; since the above librations are not correlated, the total libration of the 4:3 MMR will be large. Consequently, the total libration amplitude of $\phi_{4:3}$ will amount to $\sim 170^{\circ}$, from the libration of Pluto relative to Neptune in our simulations, plus twice the librating amplitude of any particular Twotino relative to Neptune.
@@
In principle, this would restrict the 4:3 MMR to Twotinos with librations smaller than $95^{\circ}$ (i.e. Twotinos in the leading and trailing islands); however, objects where librations of $\phi_{4:3}$ go beyond $360^{\circ}$ are not circulating in the most conventional sense, as they will be going always forward exactly once, for Twotinos whose libration relative to Neptune lies between $95^{\circ}$ and $275^{\circ}$, or twice, for Twotinos that librate between $275^{\circ}$ and $360^{\circ}$, and then they will circulate backwards the same number of turns, always being bound by up to $890^{\circ}$; therefore, from here on we will call this behavior as a MMR regardless of the amplitude of libration of $\phi_{4:3}$. The resulting resonant effect will be present (as we will prove in Section 4), though diluted, similar to the weakening present in higher-order MMRs (such as the 5:3, 7:4, etc., see e.g., Tamayo and Hadden, 2025).
Technically, the 4:3 MMR is a 6:4:3 three-body resonance between Neptune:Pluto:Twotinos, but the effects of Neptune on the Twotinos and Pluto through the 2:1 (6:3) and 3:2 (6:4) MMRs have been studied extensively (e.g., Williams and Benson, 1971; Malhotra, 1996; Gallardo and Ferraz-Mello, 1998; Nesvorný and Roig, 2001; Kotoulas, 2005), thus it is easier to consider the effects of the 4:3 resonance as an isolated two-body resonance than to examine the full scope of the 6:4:3 three-body resonance.
@@
Figure 2: Co-rotational orbits with respect to Pluto of the four sample Twotinos (top panels) over 1 Myr, as well as the measured probability distribution function to find each Twotino at a certain co-rotating angle (bottom panels). The black curves represent Pluto’s orbit in the co-rotational frame, while colored curves represent each Twotino, following the same color scheme as Figure 1. To see the evolution of the orbit, see the attached animation of these four objects on the GitHub repository.
@@
For our numerical experiments, we followed a similar procedure as in Muñoz-Gutiérrez et al. (2025). We conducted short-term, 10 Myr simulations of the 152 Twotino candidates retrieved from the JPL’s Small-Body Database  [^2] https://ssd.jpl.nasa.gov/horizons/ with a semimajor-axis in the range $47<a<48.5$ au. Our simulations include the Sun, the four giant planets, and Pluto [^3] We consider a single body with a mass equal to the mass of Pluto and Charon together, with masses taken from Stern et al. (2018). as massive bodies.
@@
We perform our simulations with REBOUND (Rein and Liu, 2012), using the MERCURIUS (Rein et al., 2019) integrator. We set an initial time-step of 180 days and an accuracy parameter of $10^{-9}$. We use a 5-yr output cadence to follow the evolution of the resonant orbits with high precision; this results in $2\times 10^{6}$ data points. Following equations 2 and 3, we obtain the 2:1 MMR resonant argument with respect to Neptune, and the 4:3 MMR resonant argument with respect to Pluto for each of the Twotinos. Overall, we obtain 91 resonant objects in the 2:1 MMR with Neptune, i.e., those with libration amplitudes lower than $340^{\circ}$ in the 10 Myr time interval of the integration. Out of these objects, 35 and 21 Twotinos librate in the leading and trailing islands, respectively, while 35 librate in the symmetric resonant island.
@@
Figure 1 shows the 2:1 and 4:3 resonant arguments over 1 Myr of four selected objects, representing the leading, trailing, and symmetric islands of the 2:1 MMR, as well as a non-resonant (circulating) object. The top panel shows the evolution of the 2:1 resonant argument, while the bottom panel corresponds to twice the 2:1 resonant argument minus Pluto’s resonant argument ($\phi_{3:2}$), as described in equation 3. It can be seen how the leading object (blue curve) closely follows Pluto’s resonant argument, as it has a near-zero libration amplitude in the 2:1 MMR leading island. In contrast, the object on the trailing island (green curve) exhibits amplitude modulations throughout its evolution, with full cycles taking close to $0.3$ Myr to complete. The object in the symmetric island (red curve) shows secondary peaks, where the slower evolution will result in preferred longitudinal angles. Finally, the non-resonant orbit (gray curve) circulates with a continuously increasing amplitude. The behavior of the gray curve in the bottom panel of Figure 1 is due to the way we defined $\phi_{4:3}$, and the fact that $\phi_{2:1}$ is discontinuous while limited to the $0^{\circ}$ - $360^{\circ}$ range.
@@
All objects in the symmetric island (as well as 5 objects in the asymmetric islands) have $\phi_{4:3}$ libration amplitudes larger than $360^{\circ}$; in fact, most of them have libration amplitudes larger than $720^{\circ}$. Although this could be interpreted as a form of circulation, this circulation is not the same as that observed in the non-resonant example. In conventional three-body systems, once $360^{\circ}$ degrees have been passed, the same configuration is reached again. In our case, the influence of the resonances with Neptune prevents the system from returning to the same position. This kind of behavior of $\phi_{4:3}$ will cause certain angles to be preferred.
@@
All remaining Twotinos show behaviors similar to the above four examples, described by the linear combination of $\phi_{2:1}$ and $\phi_{3:2}$, although with secondary amplitude modulations due to varying frequencies between the 3:2 and 2:1 resonant arguments. The complete set is uploaded to a GitHub repository along with our classification for each of the 91 Twotinos (as well as 61 circulating objects near the resonance)  [^4] https://github.com/Sramrev/Pluto-Twotino-MMR.git. This confirms that a 2:1 MMR between a TNO and Neptune directly correlates with a 4:3 MMR between the same object and Pluto.
@@ 4 Dynamical Significance of the 4:3 Resonance
To further exemplify the significance of the Twotino-Pluto MMR, Figure 2 shows the co-rotating orbits of Twotinos relative to Pluto, corrected via the mean longitude of the latter. Objects in the asymmetric islands, with libration amplitudes of $174^{\circ}$ for the leading Twotino (in blue) and $229^{\circ}$ for the trailing Twotino (in green), occupy the same longitudinal angles in the co-rotational frame at perihelion. The white curves superimposed on the colored curves represent the first 62,000 yr of each orbit, which is equivalent to the slower period of our example Twotinos, i.e., the one librating in the symmetric island (shown in red). This period is also shown by the vertical dashed lines in figure 1. The confinement of the allowed angles in a co-rotational frame to a restricted range of values is characteristic of resonant dynamics. In the case of these Twotinos, such confinement allows them to avoid, or to regulate, the strength of gravitational perturbations from Pluto during conjunctions.
@@
This behavior is further confirmed in the lower panels of Figure 2, which show the probability distribution of visited co-rotating angles for each of the selected Twotinos. In these panels, we can see how the leading and trailing objects have clear preferred angles. Among these 56 objects, the ratios between favored and unfavored angles cover a 1.5 - 4 factor range. The librating objects with $360^{\circ}<\phi_{4:3}<890^{\circ}$ show the same effect as particles with librations $\phi_{4:3}<360^{\circ}$, although the peaks are less pronounced; the sample symmetric object (shown in red) has the second highest signal among objects in the symmetric island, with a favored to unfavored angles ratio of 1.7; all other symmetric Twotinos have values between 1 and 1.9, with an average value of $\sim 1.3$.
Even when the libration amplitude of $\phi_{4:3}$ exceeds $360^{\circ}$, the resonant configuration in a co-rorational frame allows for the repetition of pulls between objects in a limited set of directions without being canceled by pulls from all other directions. This is in striking contrast to the circulating particles in the 2:1 MMR, one such example shown in gray in Figure 2, which, apart from covering almost homogeneously the entire circumference in the co-rotating frame, shows no preferred angles in the probability density histogram.

### v2

@@ Abstract
Methods. We run high-resolution, 10 Myr REBOUND simulations of the observed Twotino population in the Kuiper belt, under the perturbations by the Sun, the four giant planets, and Pluto, as massive objects.
Results. We find that all objects trapped in the 2:1 MMR with Neptune are locked in a weak 4:3 MMR with Pluto. The 4:3 resonant argument of most objects trapped in the leading and trailing islands of the 2:1 MMR, librates with amplitudes lower than $360^{\circ}$. Objects in the symmetric islands of the 2:1 MMR librate in the 4:3 MMR with amplitudes greater than $360^{\circ}$, but, contrary to circulating objects, will oscillate by up to $850^{\circ}$ visiting preferred regions on Pluto’s co-rotating frame, indicating a diluted resonant effect that may also perturb their orbits on secular timescales.
Conclusions. The importance of Pluto in shaping the structure of the trans-Neptunian region should be reconsidered, specifically for resonant populations. Moreover, with current computational power, its exclusion from simulations cannot be justified.
@@ 1 Introduction
Long after Clyde Tombaugh’s discovery of Pluto in 1930, the first long-term integration of its orbit (Cohen and Hubbard, 1965), aided by the first computers, enabled the determination of its resonant state with Neptune. Some of the pioneering works that studied the secular dynamics of the outer Solar System, from the late 1980s to the early 1990s, considered the perturbations of the Sun and the four giant planets, together with an ever larger number of test particles (see, e.g. Duncan et al., 1988; Torbett, 1989; Gladman and Duncan, 1990; Levison and Duncan, 1993). It probably seemed wasteful to include other massive bodies, since Uranus is over 6,500 times more massive than Pluto [^1] See e.g., https://ssd.jpl.nasa.gov/planets/phys_par.html, and references therein., and over 5,200 times more massive than Eris, the two most massive currently known trans-Neptunian objects (TNO).
It has also been recognized that Pluto has an important influence over objects in its vicinity, being responsible, among other things, for the destabilization of other Plutinos, which contribute to the injection rate of comets into the inner Solar System (Morbidelli, 1997; Yu and Tremaine, 1999; Nesvorný et al., 2000; Tiscareno and Malhotra, 2009; Muñoz-Gutiérrez et al., 2019).
However, Pluto can have a secular effect on other objects besides its neighboring Plutinos. In Muñoz-Gutiérrez et al. (2026, hereafter Paper I), we found that Pluto has a significant effect on the leaking rate of the 2:1 mean motion resonance (MMR); especifically, the stable fraction of Twotinos after 4 Gyr of evolution is reduced from 47% when considering only the giant planets, to 19% when including Pluto along the giant planets as a massive object in simulations.
The origin of this significant effect is not straightforwardly related to Pluto’s mass. In Paper I, we found that Eris with a similar mass, produces an effect an order of magnitude smaller than Pluto on the Twotinos. We note, however, that although Pluto is not massive enough to capture other TNOs into its own MMRs, Neptune has captured enough objects in its resonances, and those will therefore be at the same time in resonance with Pluto.
In this letter, we show the geometrical configuration between Pluto and Twotinos corresponds to a previously unrecognized 4:3 MMR between them, and that the cumulative resonant effects from Pluto will slowly but surely distort Twotinos’ orbits. This shows that Pluto is an important ingredient when studying the secular evolution of the outer Solar System, and with current computational capabilities, its exclusion from simulations cannot be justified.
Figure 1: Evolution of the 2:1 resonant arguments with respect to Neptune (top panel) and the 4:3 resonant arguments with respect to Pluto (bottom panel) for four selected objects. The blue ($2015\,{\rm VH}_{166}$, leading), green ($2015\,{\rm GZ}_{54}$, trailing), and red ($2015\,{\rm KX}_{173}$, symmetric) curves show the behavior in the three possible Twotino libration islands. The gray curve shows a circulating ($1999\,{\rm RX}_{215}$, non-resonant) object. Each Twotino has a corresponding dashed line, indicating the evolution of the same object, but without the gravitational influence of Pluto; at this timescale, differences between the two scenarios are only noticeable for the Twotino in the symmetric island. The vertical dashed line is placed at $62,000$ yr, representing one libration period of the symmetric object.
@@ 2 The Connection Through Neptune
As both Pluto and any Twotino are locked in MMR with Neptune, they are also indirectly locked in resonance with each other. While Neptune orbits six times around the Sun, Pluto will orbit four times, and a Twotino will orbit three times. Therefore, Twotinos will be in a 4:3 MMR with Pluto.
It is well known that resonant arguments of Pluto and Twotinos in Neptune’s 3:2 and 2:1 MMRs are, respectively:
@@
$\phi_{2:1}$ $=$ $2\lambda_{\rm T}-\lambda_{\rm N}-\varpi_{\rm T}$ (2)
where $\lambda$ is the mean longitude, $\varpi$ the longitude of perihelion, and the sub-indices P, N, and T stand for Pluto, Neptune, and Twotinos, respectively. To cancel the dependence on Neptune’s mean longitude, we can subtract $\phi_{3:2}$ from twice $\phi_{2:1}$.
@@
The libration of this resonant argument corresponds to Twotinos being locked in a first-order, 4:3 MMR with Pluto (with $\phi_{4:3}$ corresponding to eq. 4D1.4 in Murray and Dermott, 1999). This resonance is weaker than the ones involving Neptune, as its relative strength includes a third-order eccentricity term ($e_{P}e_{T}^{2}$), instead of first-order terms as other resonant arguments with Neptune do ($e_{P}$, and $e_{T}$, respectively); it is also slightly more complex as it involves the longitudes of perihelion of both resonant objects.
It should be noted that the 4:3 MMR is part of a 6:4:3 three-body resonance between Neptune:Pluto:Twotinos, but the effects of Neptune on the Twotinos and Pluto through the 2:1 (6:3) and 3:2 (6:4) MMRs have been studied extensively (e.g., Williams and Benson, 1971; Malhotra, 1996; Gallardo and Ferraz-Mello, 1998; Nesvorný and Roig, 2001; Kotoulas, 2005), while the effects of Pluto on the Twotinos have not been studied previously. Here, we focus on the relationship between Pluto and the Twotinos as a first-order 4:3 MMR, without forgetting that, given Pluto’s mass being four orders of magnitude smaller than Neptune’s, the effect of Pluto on the Twotinos can be considered as a perturbation to Neptune’s dominating dynamics. Thus, it is justified to study an isolated two-body resonance, even if the factor that puts those objects in resonance is an external element. Overall, we stress that $N$-body simulations with a massive Pluto are required to correctly model the secular evolution of the trans-Neptunian region.
@@
The libration of Pluto’s 3:2 resonant argument has a large amplitude (from $\sim 160^{\circ}$ to $\sim 172^{\circ}$, e.g. Milani et al., 1989; Malhotra and Ito, 2022). There are also relatively large librations of Twotinos in the three possible resonant islands of the 2:1 MMR, especially for Twotinos in the symmetric island (e.g. Figure 4 of Paper I, ); thus, there will be some libration between Pluto and the Twotinos; since the above librations are not correlated, the total libration of the 4:3 MMR will be large. Consequently, the total libration amplitude of $\phi_{4:3}$ will amount to $\sim 170^{\circ}$, from the libration of Pluto relative to Neptune, plus twice the librating amplitude of any particular Twotino relative to Neptune.
@@
In principle, this would restrict the 4:3 MMR to Twotinos with librations smaller than $95^{\circ}$ (i.e. Twotinos in the leading and trailing islands); however, objects where librations of $\phi_{4:3}$ go beyond $360^{\circ}$ are not circulating in the most conventional sense, as they will be going always forward exactly once, for Twotinos whose libration relative to Neptune lies between $95^{\circ}$ and $275^{\circ}$, or twice, for Twotinos that librate between $275^{\circ}$ and $360^{\circ}$, and then they will circulate backwards the same number of turns, always being bound by up to $890^{\circ}$. Therefore, regardless of the amplitude of libration of $\phi_{4:3}$, we will consider such particles as resonant as long as $\phi_{4:3}$ is bounded inside those limits. Despite the large libration amplitude, the resulting resonant effect will be present (as we will prove in Section 4), though diluted, similar to the weakening present in higher-order MMRs (such as the 5:3, 7:4, etc., see e.g., Tamayo and Hadden, 2025).
@@
Figure 2: Co-rotational orbits with respect to Pluto of the four sample Twotinos (top panels) over 1.5 Myr, as well as the measured probability distribution function to find each Twotino at a certain co-rotating angle (bottom panels). The black curves represent Pluto’s orbit in the co-rotational frame, while colored curves represent each Twotino, following the same color scheme as Figure 1. To see the evolution of the orbit, see the attached animation of these four objects on the GitHub repository.
@@
For our numerical experiments, we followed a similar procedure as in Paper I. We conducted short-term, 10 Myr simulations of the 152 Twotino candidates retrieved from the JPL’s Small-Body Database  [^2] https://ssd.jpl.nasa.gov/horizons/ with a semimajor-axis in the range $47<a<48.5$ au. Our simulations include the Sun, the four giant planets, and Pluto [^3] We consider a single body with a mass equal to the mass of Pluto and Charon together, with masses taken from Stern et al. (2018). as massive bodies.
@@
We perform our simulations with REBOUND (Rein and Liu, 2012), using the IAS15 (Rein and Spiegel, 2015) integrator. We set an initial time-step of 100 days and an accuracy parameter of $10^{-9}$. We use a 5-yr output cadence to follow the evolution of the resonant orbits with high precision. Following equations 2 and 3, we obtain the 2:1 MMR resonant argument with respect to Neptune, and the 4:3 MMR resonant argument with respect to Pluto for each Twotino. Overall, we obtain 91 resonant objects librating in the 2:1 MMR with Neptune, i.e., those with libration amplitudes below a set limit of $340^{\circ}$ throughout the integration. Out of these Twotinos, 35 librate in the leading island, 21 in the trailing island, and the other 35 in the symmetric island.
@@
Figure 1 shows the 2:1 and 4:3 resonant arguments over 1.5 Myr of four selected objects, representing the leading, trailing, and symmetric islands of the 2:1 MMR, as well as a non-resonant (circulating) object. The top panel shows the evolution of the 2:1 resonant argument, while the bottom panel corresponds to the evolution of the 4:3 resonant argument, as described in equation 3. In the bottom panel of Fig. 1, the leading object (blue curve) closely follows Pluto’s resonant argument, as it has a near-zero libration amplitude in the 2:1 MMR. In contrast, the object on the trailing island (green curve) exhibits amplitude modulations throughout its evolution. The object in the symmetric island (red curve) shows secondary peaks, where the slower evolution will result in preferred longitudinal angles. In contrast, the non-resonant orbit (gray curve) circulates in both panels with a continuously increasing amplitude.
To highlight the effect of Pluto on the evolution of Twotinos, we also show a red dashed line representing the evolution of $2015\,{\rm KX}_{173}$ (the symmetric object) when turning off Pluto’s influence. Pluto’s effect seems negligible within 1 Myr, and statistically insignificant at 1.5 Myr, however after 4 Gyr, it will result in completely different orbits for some Twotinos. Although imperceptible in Fig. 1, there are dashed lines for the other 3 objects, but they need approximately 10 Myr to show noticeable differences.
@@
All objects in the symmetric island (as well as 5 objects in the asymmetric islands) have $\phi_{4:3}$ libration amplitudes larger than $360^{\circ}$; in fact, most of them have libration amplitudes larger than $720^{\circ}$. Although this could be interpreted as a form of circulation, this circulation is not the same as that observed in the non-resonant example.
@@
All remaining Twotinos show behaviors similar to the above four examples, although with secondary amplitude modulations due to varying frequencies between the 3:2 and 2:1 resonant arguments. The complete figure set is uploaded to a GitHub repository along with our classification for each of the 91 Twotinos (as well as the 61 circulating objects near the resonance)  [^4] https://github.com/Sramrev/Pluto-Twotino-MMR.git. This confirms that a 2:1 MMR between a TNO and Neptune directly correlates with a 4:3 MMR between the same object and Pluto.
@@ 4 Dynamical Significance of the 4:3 Resonance
The presence of a MMR between two objects implies not only a geometrical conmensurability, but a dynamical enhanced interaction tied to the geometrical configuration. In general, this means we could expect: (A) a persistent pp\,:q\,q ratio between the orbital periods, (B) an enhanced gravitational interaction, either between the two objects, or, at least, from the more massive object affecting the lighter one, and (C) that the gravitational perturbation carves a region of phase-space where the $p$:$q$ ratio is maintained and reinforced by the resonant interaction.
The Pluto-Twotino interaction described in the previous section fulfills two out of the three characteristics of the MMR interaction (A+B), since Pluto affects the evolution of Twotinos secularly (as has been proved in Paper I, see e.g., Fig. 7); this effect is almost certainly due to the persistent 4:3 synchronous gravitational interactions between Pluto and Twotinos. Although Pluto’s small mass is incapable of carving the Twotino phase-space (point C), they are nonetheless in interaction thanks to their mutual resonant relation with Neptune.
To further exemplify the significance of the Pluto-Twotino MMR, Figure 2 shows the co-rotating orbits of Twotinos relative to Pluto, corrected via the mean longitude of the latter. Objects in the asymmetric islands, with libration amplitudes of $174^{\circ}$ for the leading Twotino (in blue) and $229^{\circ}$ for the trailing Twotino (in green), occupy the same longitudinal angles in the co-rotational frame at perihelion. The white curves superimposed on the colored curves represent the first 62,000 yr of each orbit, equivalent to the slower libration period of our example Twotinos. The confinement of the allowed angles in a co-rotational frame to a restricted range of values is characteristic of resonant dynamics. In the case of these Twotinos, such confinement allows them to avoid, or to regulate, the strength of gravitational perturbations from Pluto during conjunctions.
@@
This behavior is further confirmed in the lower panels of Figure 2, which show the probability distribution of visited co-rotating angles for each of the selected Twotinos. In these panels, the leading and trailing objects have clear preferred angles. Among the 56 objects in the asymmetric islands, the ratios between favored and unfavored angles cover a 1.5 - 4 factor range. Librating objects with $360^{\circ}<\phi_{4:3}<850^{\circ}$ show the same effect as particles with librations $\phi_{4:3}<360^{\circ}$, although the peaks in the histograms are less pronounced; the sample symmetric object has the second highest signal among objects in the symmetric island, with a favored to unfavored angles ratio of 1.7; all other symmetric Twotinos have values between 1 and 1.9, with an average value of $\sim 1.3$.
This is in striking contrast to the circulating particles near the 2:1 MMR, one such example shown in gray in Figure 2, which, apart from covering the entire circumference in the co-rotating frame, shows no preferred angles in the probability density histogram.